OpenStudy (anonymous):

An automobile radiator contains 8 liters of a solution that is 10 percent antifreeze and 90 percent water. How much of the solution should be drained and replaced with pure antifreeze to obtain a solution that is 25 percent antifreeze? Please help with complete solutions thank you :)

OpenStudy (zpupster):

x= 10% of the remaining liquid y= the 100% anti freeze we add We know the radiator contains 8L so x + y = 8 but we want to that to equal a 25 % of 8L system of equations is .1x+1y= .25(8) x + y = 8 can you do it from here?

OpenStudy (anonymous):

I'll try thank you so much @zpupster :)

OpenStudy (zpupster):

you're welcome

OpenStudy (anonymous):

@zpupster I got stuck with the .1x + 1y = .25(8)

OpenStudy (zpupster):

ok i was away .1x+1y= .25(8) x + y = 8 .1x+1y= .25(8) .25 is also 1/4 so 8*(1/4) is 2 and y = 1y we have now .1x+y =2 x+y=8 or 8-x=y substitute that in for y in .1x + (8-x) =2 then -1x+.1x + 8 = 2 combine like terms and subtract 8 both sides -.9x=-6 divide both sides by .9 what did you get

OpenStudy (zpupster):

hope i helped i have to be out for awile.

OpenStudy (anonymous):

Thank you uhh I got 6.66667

OpenStudy (zpupster):

yes (or 6 and 2/3) Since the quantities add to 8L, y = 4/3. So drain 4/3 L add 4/3 L pure antifreeze to get 25% anti freeze

OpenStudy (anonymous):

Ohhh! Thank you, thank you very much :D